OpenStudy (anonymous):

lim x->0 sin4x/tan2x

OpenStudy (agreene):

\[\lim_{x \rightarrow 0} \frac{\sin(4x)}{\tan(2x)}\] use l'hopitals: lim x->0 f(x)/g(x) = lim x->0 f'(x)/g'(x) to get: \[\lim_{x \rightarrow 0} \frac{4\cos(4x)}{2\sec^2(2x)}\] then you can just plug in 0 because it is defined, and get \[\frac{4}{2}=2\]

OpenStudy (anonymous):

Thanks, but I haven't done calculus yet so, is there another method?

OpenStudy (agreene):

Not that would be easier... actually, I can't really think of anything else you can do to evaluate it, since it is undefined at 0.

OpenStudy (anonymous):

Ok then. The topic we're doing right now is trigonometry. Do you think \[sinx/x =0 \] can be used?

OpenStudy (agreene):

I suppose you can try rearranging things and using double angle identities and hope something cancels... but that's a ton more work O.o